OpenStudy (anonymous):

Use the definition of the derivative to find f'(x) if f(x) = 2x / (x=4) Does this mean first principles?

myininaya (myininaya):

So you are finding the derivative of f(x)=2x at x=4?

myininaya (myininaya):

\[f'(x)=\lim_{h \rightarrow 0}\frac{f(x+h)-f(x)}{h}\] this definition of derivative

OpenStudy (anonymous):

no i'm just finding the derivative using first principles can you help me with that?

myininaya (myininaya):

so if f(x) isn't 2x then what is f(x)?

myininaya (myininaya):

was f(x) suppose to be 2x/(x+4) or 2x/(x-4) instead of it being 2x ?

OpenStudy (anonymous):

+4

myininaya (myininaya):

2x/(x+4) ?

OpenStudy (anonymous):

yes

myininaya (myininaya):

well if f(x)=2x/(x+4) do you know how to find f(x+h)?

OpenStudy (anonymous):

just plug it in

myininaya (myininaya):

yep where the x is

OpenStudy (anonymous):

yes but i can't get the real derivative. can you show me the steps? i know the answer is 8/(x+4)^2

myininaya (myininaya):

\[f'(x)=\lim_{h \rightarrow 0}\frac{\frac{2(x+h)}{x+h+4}-\frac{2x}{x+4}}{h}\] combine top fractions there it is easy if you just use the drunken smiley face method :)